### v1

@@ 5 Multi-dimensional Refusal Cones
We extend RDO to higher dimensions by searching for regions in activation space where all vectors control refusal behavior. For this, we optimize an orthonormal basis $\mathcal{B}=[{{\bm{b}}}_{1},\dots,{{\bm{b}}}_{N}]$ spanning an $N$-dimensional polyhedral cone $\mathcal{R}_{N}=\{\sum_{i=1}^{N}\lambda_{i}{{\bm{b}}}_{i}\;|\;\lambda_{i}\geq 0\}\backslash\{\mathbf{0}\}$, where all directions ${{\bm{r}}}\in\mathcal{R}_{N}$ satisfy the refusal properties (Definition 4.1). Since all directions in the cone correspond to the same refusal concept, we also refer to this as a concept cone. The constraint $\lambda_{i}\geq 0$ ensures that all directions within the cone consistently strengthen refusal behavior. Without this constraint, allowing negative coefficients could introduce opposing effects, reducing the overall effectiveness. Enforcing orthogonality of the basis vectors prevents finding co-linear directions. Note that in practice, directions in activation space cannot be scaled arbitrarily high without model degeneration, which effectively bounds $\lambda_{i}$.
@@
8:     $\mathcal{B}\leftarrow$ $(\mathcal{B})$
@@
In Algorithm 2, we describe the procedure to find the cone’s basis vectors. The basis vectors are initialized randomly and iteratively optimized using projected gradient descent. We compute the previous losses defined in Algorithm 1 on Monte Carlo samples from the cone, as well as on the basis vectors themselves. Computing the loss on the basis vectors improves both stability and the lower bounds of the ASR. This is because the basis vectors are the boundaries of the cone and thus tend to degrade first. After each step, we project the basis back onto the cone using the Gram-Schmidt orthogonalization procedure. Because the directional ablation operation uses the normalized ${\hat{{{\bm{r}}}}}$ rather than ${{\bm{r}}}$, sampling convex combinations of the basis vectors and normalizing them would introduce a bias towards the basis vectors themselves. Instead, we sample unit vectors in the cone uniformly to ensure better coverage of the space.
@@
Can we find refusal concept cones? We train cones of increasing dimensionality using the same experimental setup as described in Section 4. We measure the cone’s effectiveness in mediating refusal by sampling 256 vectors from each cone and computing the ASRs of the samples for directional ablation. We show the results in Figure 3 and confirm that the directions in the cones have the desired refusal properties in Figure 14. Notably, we identify refusal-mediating cones with dimensions up to five across all tested models. This suggests that the activation space in language models exhibits a general property where refusal behavior is encoded within multi-dimensional cones rather than a single linear direction.
Do larger models contain higher-dimensional cones?
In Figure 4, we evaluate the effect of model size within the Qwen 2.5 family. We observe that across all model sizes, the lower bounds of cone performance degrade significantly as dimensionality increases. In other words, a higher number of sampled directions have low ASR. Larger models appear to support higher-dimensional refusal cones. A plausible explanation is that models with larger residual stream dimensions (e.g., 5120 for the 14B model vs. 1536 for the 1.5B model) allow for more distinct and orthogonal directions that mediate refusal. Finally, in Figure 11, we confirm that directions sampled from these cones effectively induce refusal behavior, further supporting the notion that multiple axes contribute to the model’s refusal decision.
@@
To further investigate the role of different vectors, we assess whether multiple sampled cone directions influence the model in complementary ways. Specifically, we sample varying numbers of directions from Gemma-2-2B’s four-dimensional refusal cone and, for each prompt, select the most effective one under directional ablation (more details in Appendix A). To ensure a fair comparison, we use temperature sampling with the single-dimension RDO direction to generate the same number of attacks and similarly select the most effective instance. We study Gemma 2 2B and sample from its four-dimensional cone, since performance degrades significantly for larger dimensions (see Figure 10).
Figure 5 shows that sampling multiple directions leads to higher ASR compared to sampling with various temperatures in the low-sample regime. For a higher number of samples, the randomness dominates the success of the attack. However, the higher ASR in the low-sample regime suggests that different directions capture distinct, complementary aspects of the refusal mechanism. Additionally, Figure 13 reveals that ASR increases with cone dimensionality but plateaus at four dimensions. This trend indicates that higher-dimensional cones offer an advantage over single-direction manipulation, likely by influencing complementary mechanisms. The plateau likely occurs because the model does not support higher-dimensional refusal cones.
Figure 5: ASR for best-of-N sampling using $N$ samples from the 4-dimensional refusal cone of Gemma-2-2B, compared to best-of-N sampling with temperature $T$ using the single-dimension RDO.
Key Takeaways. We show that refusal mechanisms in LLMs span high-dimensional polyhedral cones, capturing diverse aspects of refusal behavior. This highlights their geometric complexity and demonstrates the effectiveness of our gradient-based method in identifying intricate structures.
@@ 6 Mechanistic Understanding of Directions
In the previous section, we demonstrated that refusal behavior spans a multi-dimensional cone with infinitely many directions. However, whether the orthogonal refusal-mediating basis vectors manipulate independent mechanisms remains an open question. In this section, we conduct a mechanistic analysis to investigate how these directions interact within the model’s activation space and whether they can be directly influenced through input manipulation. This allows us to determine whether they are merely latent properties of the network or actively utilized by the model in response to specific prompts.
@@ 6.1 Representational Independence
However, despite this mathematical property, recent work by Park et al. (2024) suggests that in language models, conclusions about causal independence cannot be drawn using orthogonality measured with the Euclidean scalar product. Although their assumptions differ from ours, especially since they assume a one-to-one mapping from output feature to direction in activation space, their experiments suggest that independent directions are almost orthogonal. This motivates a deeper empirical examination of how orthogonal refusal directions in language models interact in practice.
Are orthogonal directions independent? To explore this, we first use RDO to identify a direction ${{\bm{r}}}$ for Gemma 2 2B that is orthogonal to the DIM direction ${{\bm{v}}}$, i.e., ${{\bm{r}}}^{\top}{{\bm{v}}}=0$. We then measure how much one direction is influenced when ablating the other direction by monitoring the cosine similarity $\smash{\cos(\lambda,\mu)=\frac{\lambda^{\top}\mu}{\lvert\lvert\lambda\rvert\rvert\cdot\lvert\lvert\mu\rvert\rvert}}$ between the prompt’s representation in the residual stream ${\bm{x}}$ and the directions ${{\bm{v}}}$ and ${{\bm{r}}}$. Specifically, we track: $\smash{\cos({{\bm{r}}},{\bm{x}}_{i}^{(l)}({{p}_{\text{harm}}}))}$ and $\smash{\cos({{\bm{v}}},{\bm{x}}_{i}^{(l)}({{p}_{\text{harm}}}))}$ at the last token position and for all layers $\smash{l\in\{0,\dots,L\}}$ on 128 harmful instructions in our validation set. Intuitively, ablating a causally independent direction in earlier layers should not intervene with the reference direction in later layers. Otherwise, there is some indirect influence through the non-linear transformations of the neural network.
The top row of Figure 6 shows how the cosine similarity between the RDO and DIM directions changes under intervention. The left plot shows the cosine similarity between the RDO direction and the activations on a normal forward pass (solid line) and while ablating the DIM direction (dashed line). The right plot presents the reverse setting. Despite enforced orthogonality, ablating RDO indirectly reduces the representation of the DIM direction in the model activations in the later layers, as measured by cosine similarity. This effect is reciprocal, suggesting that orthogonality alone does not guarantee independence throughout the network.
@@
$\begin{split}\forall l\in L:\cos({\bm{x}}^{(l)},\lambda)&=\cos(\tilde{{\bm{x}}}^{(l)}_{abl(\mu)},\lambda)\\ \text{and}\;\cos({\bm{x}}^{(l)},\mu)&=\cos(\tilde{{\bm{x}}}^{(l)}_{abl(\lambda)},\mu).\end{split}$
This means that, instead of relying solely on orthogonality, we define two directions as representationally independent if ablating one has no effect on how much the other is represented in the model activations. To enforce this property, we extend Algorithm 1 with an additional loss term that penalizes changes in cosine similarity at the last token position when ablating on harmful instructions:
$\begin{split}\mathcal{L}_{\text{RepInd}}=\frac{1}{|L|}\sum_{l\in L}\Big{[}\big{(}\cos(x^{(l)},{{\bm{r}}})&-\cos(\tilde{x}^{(l)}_{abl({{\bm{v}}})},{{\bm{r}}})\big{)}^{2}\\ +\big{(}\cos(x^{(l)},{{\bm{v}}})&-\cos(\tilde{x}^{(l)}_{abl({{\bm{r}}})},{{\bm{v}}})\big{)}^{2}\Big{]}.\end{split}$
Figure 7: Performance of RepInd directions. Each direction is representationally independent to all previous directions and the DIM direction.
Do independent directions exist? With this extension, we can find a direction that is RepInd from the DIM direction, yet still fulfills the refusal properties from Definition 4.1. We show the representational independence in the second row of Figure 6, where we see that the RepInd and DIM direction barely affect each other’s representation under directional ablation.
We iteratively search for additional directions that are not only RepInd to DIM but also of all previously identified RepInd directions. Despite these strong constraints, we successfully identify at least five such directions that maintain an ASR significantly above random vector intervention (Figure 7), as well as a refusal cone with RepInd basis vectors (Figure 12). However, in Figure 7 and Figure 12 performance degrades more rapidly compared to the results in Section 4 and Section 5. This decline could be attributed to the increased difficulty of the optimization problem due to additional constraints. Alternatively, it may suggest that Gemma 2 2B possesses a limited number of directions that independently contribute to refusal. If the latter is true, this implies that the directions in the refusal cones exhibit non-linear dependencies. Nevertheless, these results show that refusal in LLMs is mediated by multiple independent mechanisms, underpinning the idea that refusal behavior is more nuanced than previously assumed.
@@ 6.2 Manipulation from input
Can we access these directions from the input? Having found several independent directions that are distinct from DIM, we investigate whether these directions can ever be ”used” by the model, by checking if they are accessible from the input or if they live in regions that no combination of input tokens activates. To this end, we use GCG Zou et al. (2023b) to train adversarial suffixes, which are extensions to the prompts that aim to circumvent the safety alignment. In addition to the standard cross-entropy loss on an affirmative target, we add a loss term that incentivizes the suffix to ablate RepInd 1.
Figure 8: Representation of RepInd 1 in model activations on harmful instructions before and after adversarial attacks with GCG.
In Figure 8, we show the cosine similarities between RepInd 1 and the model activations on both harmful prompts ${{p}_{\text{harm}}}$ from JailbreakBench and the same prompts with adversarial suffixes ${p}_{\text{adv}}$. We observe that GCG is able to create suffixes that significantly reduce how much RepInd 1 is represented. These suffixes successfully jailbreak the model $36\%$ of the time, which is similar to the ASR of RepInd 1.
@@ 7 Limitations
While our work provides new insights into the geometry of refusal in LLMs, some limitations remain. The refusal directions we compute are all optimized on the same targets, which may limit their ability to capture fully distinct mechanisms. Extending our method to incorporate diverse targets or leveraging reinforcement learning with a judge-based reward function could help identify additional independent mechanisms Geisler et al. (2025). Furthermore, while we establish the existence of higher-dimensional refusal cones, we cannot rule out the possibility of other yet-undiscovered regions in the model that mediate refusal.
@@ 8 Conclusion
This work advances the understanding of refusal mechanisms in LLMs by introducing gradient-based representation engineering as a powerful tool for identifying and analyzing refusal directions. Our method yields more effective refusal directions with fewer side effects, demonstrating its viability for extracting meaningful structures while allowing for greater modeling flexibility. We establish that refusal behaviors can be better understood via high-dimensional polyhedral cones in activation space rather than a single linear direction, highlighting their complex spatial structures. Additionally, we introduce representational independence and show that within this space of independent directions multiple refusal directions exist and correspond to distinct mechanisms. Our gradient-based representation engineering approach can be extended to identify various concepts beyond refusal by simply changing the optimization targets. The generated findings provide new insights into the geometry of aligned LLMs, highlighting the importance of structured, gradient-based approaches in LLM interpretability and safety.
@@ A.1 Datasets
We construct our experimental dataset using harmful and harmless instructions from established benchmarks. For harmful instructions, we draw from SALADBench Li et al. (2024a), a comprehensive collection of adversarial prompts from diverse sources. We exclude the Multilingual Wang et al. (2023) and ToxicChat Lin et al. (2023) sources since they are unsuited as harmful instructions. Afterwards, we sample up to 256 instructions from each remaining source. This results in 1,184 instructions for training and 128 for validation. We sample equal numbers of harmless instructions from the Alpaca dataset, and additionally reserve 128 more harmless instructions for testing.
@@ A.3 Hyperparameters and Implementation
Implementation and Evaluation Framework. All algorithms and exploratory experiments are implemented using the NNsight Fiotto-Kaufman et al. (2024) library. Additionally, we use the LM Evaluation Harness Gao et al. (2024) to run TruthfulQA Lin et al. (2021) with default settings, with the exception that we enable the use of each model’s default chat templates.
@@
where $\mathcal{I}$ contains the target token indexes and the last instruction token’s index, the subscript ${i,t}$ denotes the model output at sequence position $i$ and vocabulary index $t$ as defined in Section 2, and $+$ denotes concatenation.
@@
Appendix B Additional Experiments
In this section, we present further experimental results. Figure 9 demonstrates that adding the refusal direction successfully induces refusal behavior across all models for both DIM and RDO. Similarly, Figure 10 illustrates the refusal cone performance for Gemma 2, confirming the existence of higher-dimensional refusal cones within the Gemma 2 family. The results suggest that the maximum cone dimensionality may be four, as the lower bounds of the ASR drop sharply beyond this point. In Figure 11 we apply refusal cones to various Qwen 2.5 models across different dimensions, revealing that inducing refusal is significantly easier than conducting an attack. Notably, most directions even in high-dimensional cones remain effective at inducing refusal responses.

### v2

@@ 5 Multi-dimensional Refusal Cones
[…]
@@
8:  $\mathcal{B}\leftarrow$ GramSchmidt()($\mathcal{B}$)
@@
In Algorithm 2, we describe the procedure to find the cone’s basis vectors. The basis vectors are initialized randomly and iteratively optimized using projected gradient descent. We compute the previous losses defined in Algorithm 1 on Monte Carlo samples from the cone, as well as on the basis vectors themselves. Computing the loss on the basis vectors improves both stability and the lower bounds of the ASR. This is because the basis vectors are the boundaries of the cone and thus tend to degrade first. After each step, we project the basis back onto the cone using the Gram–Schmidt orthogonalization procedure. Because the directional ablation operation uses the normalized $^{\hat{{{\bm{r}}}}}$ rather than $\bm{r}$}}, sampling convex combinations of the basis vectors and normalizing them would introduce a bias towards the basis vectors themselves. Instead, we sample unit vectors in the cone uniformly to ensure better coverage of the space.
@@
Can we find refusal concept cones? We train cones of increasing dimensionality using the same experimental setup as described in Section 4. We measure the cone’s effectiveness in mediating refusal by sampling 256 vectors from each cone and computing the ASRs of the samples for directional ablation. We show the results in Figure 3 and confirm that the directions in the cones have the desired refusal properties in Figure 20. Notably, we identify refusal–mediating cones with dimensions up to five across all tested models. This suggests that the activation space in language models exhibits a general property where refusal behavior is encoded within multi–dimensional cones rather than a single linear direction.
Do larger models contain higher–dimensional cones?
In Figure 4, we evaluate the effect of model size within the Qwen 2.5 family. We observe that across all model sizes, the lower bounds of cone performance degrade significantly as dimensionality increases. In other words, a higher number of sampled directions have low ASR. Larger models appear to support higher–dimensional refusal cones. A plausible explanation is that models with larger residual stream dimensions (e.g., 5120 for the 14B model vs. 1536 for the 1.5B model) allow for more distinct and orthogonal directions that mediate refusal. Finally, in Figure 18, we confirm that directions sampled from these cones effectively induce refusal behavior, further supporting the notion that multiple axes contribute to the model’s refusal decision.
@@
To further investigate the role of different vectors, we assess whether multiple sampled cone directions influence the model in complementary ways. Specifically, we sample varying numbers of directions from Gemma–2–2B’s four–dimensional refusal cone and, for each prompt, select the most effective one under directional ablation (more details in Appendix A). To ensure a fair comparison, we use temperature sampling with the single–dimension RDO direction to generate the same number of attacks and similarly select the most effective instance. We study Gemma 2 2B and sample from its four–dimensional cone, since performance degrades significantly for larger dimensions (see Figure 17).
Figure 5 shows that sampling multiple directions leads to higher ASR compared to sampling with various temperatures in the low–sample regime. For a higher number of samples, the randomness dominates the success of the attack. However, the higher ASR in the low–sample regime suggests that different directions capture distinct, complementary aspects of the refusal mechanism. Additionally, Figure 19 reveals that ASR increases with cone dimensionality but plateaus at four dimensions. This trend indicates that higher–dimensional cones offer an advantage over single–direction manipulation, likely by influencing complementary mechanisms. The plateau likely occurs because the model does not support higher–dimensional refusal cones.
Figure 5: ASR for best–of–N sampling using NN samples from the 4–dimensional refusal cone of Gemma–2–2B, compared to best–of–N sampling with temperature $T$ using the single–dimension RDO.
Key Takeaways. We show that refusal mechanisms in LLMs span high–dimensional polyhedral cones, capturing diverse aspects of refusal behavior. This highlights their geometric complexity and demonstrates the effectiveness of our gradient–based method in identifying intricate structures.
@@ 6 Mechanistic Understanding of Directions
In the previous section, we demonstrated that refusal behavior spans a multi–dimensional cone with infinitely many directions. However, whether the orthogonal refusal–mediating basis vectors manipulate independent mechanisms remains an open question. In this section, we conduct a mechanistic analysis to investigate how these directions interact within the model’s activation space and whether they can be directly influenced through input manipulation. This allows us to determine whether they are merely latent properties of the network or actively utilized by the model in response to specific prompts.
@@ 6.1 Representational Independence
However, despite this mathematical property, recent work by Park et al. (2024) suggests that in language models, conclusions about causal independence cannot be drawn using orthogonality measured with the Euclidean scalar product. Although their assumptions differ from ours, especially since they assume a one–to–one mapping from output feature to direction in activation space, their experiments suggest that independent directions are almost orthogonal. This motivates a deeper empirical examination of how orthogonal refusal directions in language models interact in practice.
Are orthogonal directions independent? To explore this, we first use RDO to identify a direction $\bm{r}$}} that is orthogonal to the DIM direction $\bm{v}$}}, i.e., ${{\bm{r}}}^{\top}{{\bm{v}}}=0$. We then measure how much one direction is influenced when ablating the other direction by monitoring the cosine similarity $\smash{\cos(\lambda,\mu)=\frac{\lambda^{\top}\mu}{\lvert\lvert\lambda\rvert\rvert\cdot\lvert\lvert\mu\rvert\rvert}}$ between the prompt’s representation in the residual stream $\bm{x}$} and the directions $\bm{v}$}} and $\bm{r}$}}. Specifically, we track: $\smash{\cos({{\bm{r}}},{\bm{x}}_{i}^{(l)}({{p}_{\text{harm}}}))}$ and $\smash{\cos({{\bm{v}}},{\bm{x}}_{i}^{(l)}({{p}_{\text{harm}}}))}$ at the last token position and for all layers $\smash{l\in\{0,\dots,L\}}$ on 128 harmful instructions in our validation set. Intuitively, ablating a causally independent direction in earlier layers should not intervene with the reference direction in later layers. Otherwise, there is some indirect influence through the non–linear transformations of the neural network.
The top row of Figure 6 shows how the cosine similarity between the RDO and DIM directions changes under intervention. The left plot shows the cosine similarity between the RDO direction and the activations on a normal forward pass (solid line) and while ablating the DIM direction (dashed line). The right plot presents the reverse setting. Despite enforced orthogonality, ablating RDO indirectly reduces the representation of the DIM direction in the model activations in the later layers, as measured by cosine similarity. This effect is reciprocal, suggesting that orthogonality alone does not guarantee independence throughout the network. In Figure 21 we show the results for additional models.
@@
$\begin{split}\forall l\in L:\cos({\bm{x}}^{(l)},\lambda)&=\cos(\hat{\tilde{{\bm{x}}}}^{(l)}_{abl(\mu)},\lambda)\\ \text{and}\;\cos({\bm{x}}^{(l)},\mu)&=\cos(\hat{\tilde{{\bm{x}}}}^{(l)}_{abl(\lambda)},\mu),\end{split}$
where $\hat{\tilde{{\bm{x}}}}^{(l)}_{abl(\lambda)}=\left(f^{(l)}(\hat{\tilde{{\bm{x}}}}^{(l-1)}_{abl(\lambda)})+\hat{\tilde{{\bm{x}}}}^{(l-1)}_{abl(\lambda)}\right)_{abl(\lambda)}$ denotes the activations at layer $l$ produced from the previous layer’s already ablated activations with the ablation applied again after the residual addition.
Instead of relying solely on geometric orthogonality, we say that two directions are representationally independent when ablating one of them does not change how strongly the other is expressed in the model’s activations. Because we track cosine similarity at every layer, any non–linear distortions introduced earlier in the network are captured downstream. Consequently, representational independence guarantees that—measured by cosine similarity—no linear, non–linear, or cumulative interaction in the network increases or decreases how much the other examined direction is represented.
To enforce this property, we extend Algorithm 1 with an additional loss term that penalizes changes in cosine similarity at the last token position when ablating on harmful instructions:
$\begin{split}\mathcal{L}_{\text{RepInd}}=\frac{1}{|L|}\sum_{l\in L}\Big[\big(\cos({\bm{x}}^{(l)},{{\bm{r}}})&-\cos(\hat{\tilde{{\bm{x}}}}^{(l)}_{abl({{\bm{v}}})},{{\bm{r}}})\big)^{2}\\ +\big(\cos({\bm{x}}^{(l)},{{\bm{v}}})&-\cos(\hat{\tilde{{\bm{x}}}}^{(l)}_{abl({{\bm{r}}})},{{\bm{v}}})\big)^{2}\Big].\end{split}$
Figure 7: Attack success rate for jailbreaking the model with directional ablation of representationally independent refusal directions for different models on JailbreakBench. Each direction is representationally independent to all previous directions and the DIM direction.
Do independent directions exist? With this extension, we can find a direction that is RepInd from the DIM direction, yet still fulfills the refusal properties from Definition 4.1. We illustrate the representational independence for Gemma 2 2B in the second row of Figure 6, where we see that the RepInd and DIM direction barely affect each other’s representation under directional ablation.
We iteratively search for additional directions that are not only RepInd to DIM but also of all previously identified RepInd directions. Despite these strong constraints, we successfully identify multiple such directions that maintain an ASR significantly above random vector intervention (Figure 7). The ASR declines as you search for more directions, which could be attributed to the increased difficulty of the optimization problem due to additional constraints, or that the models contain a limited number of directions that independently contribute to refusal. Nevertheless, these results show that refusal in LLMs is mediated by multiple independent mechanisms, underpinning the idea that refusal behavior is more nuanced than previously assumed.
[…]
Figure 8: Compositionally ablating the top–k RepInd directions compared to ablating the DIM direction
@@ 6.2 Manipulation from input
Can we access these directions from the input? Having found several independent directions that are distinct from DIM, we investigate whether these directions can ever be “used” by the model, by checking if they are accessible from the input or if they live in regions that no combination of input tokens activates. To this end, we use GCG (Zou et al., 2023b) to train adversarial suffixes, which are extensions to the prompts that aim to circumvent the safety alignment. In addition to the standard cross–entropy loss on an affirmative target, we add a loss term that incentivizes the suffix to ablate RepInd–1.
Figure 9: Representation of the RepInd–1 direction in model activations on harmful instructions before and after adversarial attacks with GCG.
In Figure 9, we show the cosine similarities between RepInd–1 and the model activations on both harmful prompts ${p}_{\text{harm}}$} from JailbreakBench and the same prompts with adversarial suffixes ${p}_{\text{adv}}$. We observe that GCG is able to create suffixes that significantly reduce how much RepInd–1 is represented. These suffixes successfully jailbreak the model $36\%$ of the time, which is similar to the ASR of RepInd–1.
@@ 7 Limitations
While our work provides new insights into the geometry of refusal in LLMs, some limitations remain. The refusal directions we compute are all optimized on the same targets, which may limit their ability to capture fully distinct mechanisms. Extending our method to incorporate diverse targets or leveraging reinforcement learning with a judge-based reward function could help identify additional independent mechanisms (Geisler et al., 2025). Furthermore, while we establish the existence of higher-dimensional refusal cones, we cannot rule out the possibility of other yet-undiscovered regions in the model that mediate refusal.
@@ 8 Conclusion
This work advances the understanding of refusal mechanisms in LLMs by introducing gradient-based representation engineering as a powerful tool for identifying and analyzing refusal directions. Our method yields more effective refusal directions with fewer side effects, demonstrating its viability for extracting meaningful structures while allowing for greater modeling flexibility. We establish that refusal behaviors can be better understood via high-dimensional polyhedral cones in activation space rather than a single linear direction, highlighting their complex spatial structures. Additionally, we introduce representational independence and show that within this space of independent directions multiple refusal directions exist and correspond to distinct mechanisms. Our gradient-based representation engineering approach can be extended to identify various concepts beyond refusal simply by changing the optimization targets. The generated findings provide new insights into the geometry of aligned LLMs, highlighting the importance of structured, gradient-based approaches in LLM interpretability and safety.
@@ A.1 Datasets
We construct our experimental dataset using harmful and harmless instructions from established benchmarks. For harmful instructions, we draw from SALADBench (Li et al., 2024a), a comprehensive collection of adversarial prompts from diverse sources. We exclude the Multilingual (Wang et al., 2023) and ToxicChat (Lin et al., 2023) sources since they are unsuited as harmful instructions. Afterwards, we sample up to 256 instructions from each remaining source. This results in 1,184 instructions for training and 128 for validation. We sample equal numbers of harmless instructions from the Alpaca dataset, and additionally reserve 128 more harmless instructions for testing.
@@ A.3 Hyperparameters and Implementation
Implementation and Evaluation Framework. All algorithms and exploratory experiments are implemented using the NNsight (Fiotto-Kaufman et al., 2024) library. Additionally, we use the LM Evaluation Harness (Gao et al., 2024) to run TruthfulQA (Lin et al., 2021) with default settings, with the exception that we enable the use of each model’s default chat templates.
@@
where $\mathcal{I}$ contains the target token indexes and the last instruction token’s index, the subscript i,t{i,t} denotes the model output at sequence position $i$ and vocabulary index $t$ as defined in Section 2, and ++ denotes concatenation.
@@
Appendix B Extended Results for Refusal Direction Optimization
In this section, we present additional results, including results for activation addition, more datasets for directional ablation, and more benchmarks.
B.1 Activation Addition
We first confirm that our directions can also be used to induce refusal. Figure 10 demonstrates that using RDO refusal directions for activation addition successfully induces refusal behavior across all models for both DIM and RDO, and RDO slightly outperforms DIM for most models.